Accelerator Update for June 11, 2008 - June 13, 2008
Wednesday June 11
The day shift began with Operations monitoring Tevatron (TeV) store 6214 (51.12E30) and the Recycler (RR) stash (228.6E10), with the Antiproton Source down and in access to investigate a target problem, with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.
Operations prepared the Pbar Prevault area for access at 9:56 AM.
A MI expert began a study period at 10:02 AM.
TeV experts began a D0 Interaction Point (IP) study around 12:40 PM.
The MI study ended at 1:04 PM.
TeV experts reported that they had completed the D0 IP move at 1:31 PM.
Operations terminated store 6214 at 2:27 PM.
Operations put the TeV into shot setup at 2:38 PM.
Pbar came out of access at 2:42 PM. The target was pulled out.
At 3:32 PM, a Linac RF station (LRF5) tripped off. It reset.
Operations reported at 4:46 PM that they had received the “Operational Readiness Clearance” (ORC) for the new Meson MTest experiment (T979). The old experiment, T971, will continue to run parasitically.
Operations established store 6215 at 4:58 PM with an initial luminosity of 216.03E30.
Pbar came out of access at 5:24 PM. They installed a different target.
Operations established beam to MTest at 5:25 PM.
Pbar resumed stacking at 5:35 PM.
Operations disabled the MTest beam at 6:02 PM.
At 6:08 PM, Operations reported a problem with an Accumulator RF station (ARF1). Operators contacted an expert.
At 8:06 PM, Operations prepared Pbar for an access. The expert reconnected the momentum cooling and came out of access 10 minutes later.
Pbar resumed stacking at 8:29 PM.
At 10:57 PM, LRF5 tripped off. It reset.
At 11:59 PM, Operations reported problems with Booster beam positions after transition. Operations lowered the Booster beam intensity and tuned.
Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.
Thursday June 12
The midnight shift began with Operations monitoring store 6215 (87.03E30) and the stash (73.6E10), with Pbar stacking (13.7E10), and with BNB taking beam.
At 12:28 AM, Operations reported that after the last stash a Pbar Debuncher kicker (D:EKIK) would not come on. Operators investigated and had the kicker on five minutes later. Unfortunately, the kicker tripped off two more times. At 1:03 AM, Operations put Pbar into standby and called in an expert.
The Pbar kicker expert managed to reset D:EKIK at 1:55 AM, but then it tripped off again 10 minutes later. The expert called in for more help.
At 2:42 AM, BRF8 tripped off for the first of three times in two hours. It reset.
At 3:02 AM, Operations reported that the 750KeV line power supplies began tripping off. Operators investigated and found that the power supplies were very warm to the touch. There was a fan running nearby; operators set it to a higher speed.
At 3:11 AM, LRF2 tripped off. It reset.
Operations reported at 4:40 AM that the 750 KeV line power supplies were still tripping off once or twice an hour.
The Pbar kicker experts reported at 5:01 AM that they had changed the power supply for D:EKIK.
Around 6 AM, Operations established beam to MTest, but it tripped off within a few minutes. Operators investigated and discovered that a Meson SEM (F:MW1SEM) had stopped working.
An instrumentation expert reported at 6:57 AM that the MW1SEM was broken and would have to be replaced.
Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.
At 8:03 AM, Operations reported that D:EKIK had stopped pulsing. It reset.
Operations prepared Meson enclosure C and M01 for access to replace MW1SEM at 8:37 AM.
Operations prepared MI-31 for access at 9:04 AM. They were out of access 40 minutes later.
An instrumentation expert reported at 10:04 AM that the new SEM was installed and seemed to be working properly.
Operations reestablished beam to MTest at 11:14 AM.
At 11:25 AM, a Pbar expert reported that DRF1-3 was off. He turned it on.
A TeV expert reported at 12:09 PM that he was going to move D0’s Interaction Point.
At 1:32 PM, MI suffered an abort of the MI-10 kickers (I:KPS1A, B, & C). An expert investigated.
Operations put Pbar into standby at 1:33 PM.
Linac experts began changing the driver for LRF5 at 2:18 PM.
MI experts reported at 2:20 PM that the MI-10 kicker signals were lost. The signals for the kickers come from the Booster LLRF room. Controls experts investigated and found the room very hot. The RF ENI amplifier had tripped off on an over temperature indication.
Linac experts reported that they had completed their work on LRF5. Operations established beam to BNB at 3:49 PM.
At 3:56 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to also trip. Everything reset.
Operations reported at 4:04 PM that the Recycler had stashed three times during the day shift.
At 5:58 PM, Operations reported that the Booster low level RF room was still having air conditioning problems. The Duty Mechanic put a second portable AC unit in the room and operations also placed a large fan in the room to circulate the air.
Operations disabled the MTest beam at 6:07 PM.
At 6:57 PM, LRF2 tripped off. It would not reset. Experts investigated and found a lose cable. They had the station back on by 7:29 PM.
At 7:49 PM, LRF2 began tripping off again. An expert tuned the station.
Operations reported at 8:19 PM that the temperature in the Booster LLRF room had finally cooled down enough to resume stacking.
Operations terminated store 6215 at 9:17 PM.
Operations put the TeV into shot setup at 9:57 PM.
At 10:20 PM, Operations reported that the Booster LCW temperature had alarmed on a high temperature indication and the temperature was continuing to rise. Operations contacted experts. A CUB expert was called in to bring a different heat exchanger online.
Operations established store 6216 at 10:59 PM with an initial luminosity of 191.86E30.
At 11:43 PM, Operations reported that a TeV sector A3 AC unit had tripped off and would not reset. The Duty Mechanic investigated.
At 11:51 PM, the MI-10 injection kickers tripped off. Operations called in kicker experts.
Operations reported at midnight that the Recycler had stashed two times during the evening shift.
Friday June 13
The midnight shift began with Operations monitoring store 6216 (152.88E30) and the stash (44.2E10), with Pbar stacking (13.0E10), and with BNB taking beam.
At 12:05 AM, Operations reported that a missing BEXS bit on the programmable pulse delay module was holding off the ENI amplifier.
At 12:16 AM, LRF2 tripped off. It reset.
The MI kicker experts reported at 1:31 AM that they found a lose connection on the pulse delay NIM module.
Pbar resumed stacking at 1:32 AM.
At 2:31 AM, LRF1 tripped off. It reset.
At 4:25 AM, Operations reported that the Booster LLRF room temperature was rising. The 2 nd portable AC unit was just blowing air.
Operations established beam to MTest at 6 AM.
Operations reported at 7:56 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and for the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 06/13/2008 email Fermilab|
|Security, Privacy, Legal|